# On the Element-Wise Representation and Reasoning in Zero-Shot Image Recognition: A Systematic Survey

### v2

@@ III-D Attribute Component Analysis
Attributes are the only bridge between seen and unseen classes, and they are also important supervisory signals for learning fine-grained visual representations. Therefore, the quality of the attribute space heavily determines the upper limit of the performance of a recognition system. However, most of the attributes are usually created manually or pretrained without special adjustment or filtering. Attributes contain a lot of redundant information and noise, and the complex relationships between attributes are not sorted out. As a result, the learned attribute space is not optimal. Attribute component analysis attempts to improve the above situation by adjusting or designing the attribute space. The main approaches include attribute enhancement and attribute refinement.
III-D1 Attribute Enhancement
@@
The purpose of attribute enhancement is to optimize attribute embeddings to boost the discriminability of attributes, and the main techniques include key attribute localization and attribute space optimization. The motivation of key attribute localization is similar to that of visual attention, which aids in screening by identifying and localizing the most critical parts. The cross attention mechanism mentioned in Equation 6 is one of the attribute localization techniques. In addition, LFGAA [89] designs a latent guided attention module. It assumes that each image has different attention for attributes and therefore introduces multiple subnetworks to generate exclusive attribute weights for each image. Attribute space optimization is to adjust the pretrained attribute embeddings. The original attribute embeddings have semantic overlaps affecting the discrimination, especially for fine-grained datasets. For example, a large number of attribute features are shared between different bird species, which only have minor visual variations in some of the attributes. And the original attribute embeddings are difficult to capture such attribute differences resulting in a sub-optimal space. For this reason, AREES [75] delineates the attribute embeddings with similar semantic distances into a cluster by K-means clustering algorithm, and the attributes in each cluster are subtracted from the cluster’s center of mass, i.e., the mean vector, to obtain the enhanced attribute vector. APNet [90] introduces graph modeling to discover the logical relationships between classes, and updates the attribute vectors by propagating the attribute features of neighboring nodes.
III-D2 Attribute Refinement
@@
Most algorithms just learn attribute annotations as normal labels or optimize the attribute embedding space. However, the potential irrationality of attribute annotations has not been thoroughly investigated. Since most attributes are annotated by humans, the gain of description from human perspective for deep learning is unknowable. Therefore, attribute refinement attempts to provide insights from the perspective of attribute annotation design. However, there are fewer related studies. MCZSL [91] points out that the quality of manual annotations is uncontrollable and time-consuming. Therefore, it attempts to design an automated attribute annotation scheme. Specifically, it obtains attribute descriptions from online media such as Wikipedia and designs a series of filtering strategies to filter out inappropriate attributes. DEDN [92] proposes a divide-and-conquer attribute learning scheme. It argues that attribute annotations are confusing and lead to semantic asymmetry problems. For example, for the attributes describing scenes, {studying, swimming} and {stone, tree} are clearly not in the same dimension, with the former denoting function and the latter denoting object. Therefore, it designs a dual expert network in which one expert network focuses on processing attribute information in different dimensions.
IV Zero-Shot Compositional Recognition
IV-A Problem Formulation
@@
Suppose $\mathcal{S}=\{s_{1},s_{2},...,s_{n}\}$ is a set containing $n$ state elements and $\mathcal{O}=\{o_{1},o_{2},...,o_{m}\}$ is a set containing $m$ object elements. Then a compositional set can be obtained as $\mathcal{Y}=\{(s_{1},o_{1}),(s_{1},o_{2}),...,(s_{n},o_{m})|s\in\mathcal{S},o\in\mathcal{O}\}$, which containing $n\times m$ compositions. Now we have a seen domain $D^{s}=\{(x^{s},y^{s})|x^{s}\in\mathcal{X}^{s},y^{s}\in\mathcal{Y}^{s}\}$, where $\mathcal{X}^{s}$ denotes the training image set and $\mathcal{Y}^{s}\subseteq\mathcal{Y}$ is the label set. Similarly, we have an unseen domain denoted by $D^{u}=\{(x^{u},y^{u})|x^{u}\in\mathcal{X}^{u},y^{u}\in\mathcal{Y}^{u}\}$, where $\mathcal{X}^{u}$ denotes the test image set and $\mathcal{Y}^{u}\subseteq\mathcal{Y}$ is the label set. Data in the seen and unseen domains are disjoint, i.e., $X^{s}\cap X^{u}=\phi$ and $Y^{s}\cap Y^{u}=\phi$. The objective of ZSCR is to generalize to the unseen domain by learning the element-wise knowledge in the seen domain. Specifically, for closed-world compositional recognition, the goal is to learn a mapping function $\Psi:\mathcal{X}\to\mathcal{Y}^{s}\cup\mathcal{Y}^{u}$, wherein $\mathcal{X}^{u}$ is strictly inaccessible for training. Typically, the size of $\mathcal{Y}^{s}\cup\mathcal{Y}^{u}$ is smaller than that of $\mathcal{Y}$ since some compositional are not feasible. For open-world compositional recognition [28], the goal is to learn a mapping function $\Psi:\mathcal{X}\to\mathcal{Y}$, assuming that which compositions are infeasible is unknown.
@@
We divide the cutting-edge techniques of ZSCR into three categories based on their core components, including Decoupling Representation, Dependency Modeling, and Relationship Modeling. We then further split the subcategories based on specific technical details or targets of action.
IV-B Decoupling Representation
@@
A common baseline approach is to map the image to a deep semantic space and then compute the similarity to the text embeddings and optimize with cross entropy loss. Briefly, the goal is to learn the posteriori probability $p(y|\psi(x))$, where $y=(s,o)$ and $\psi$ is a function mapping the image into the semantic space. This approach takes into account the mapping relation of compositional embeddings, yet lacks the ability to deconstruct and reorganize the visual features, thus making it difficult to generalize to the unseen domain. In other words, during training phase, the model will seek the simplest path for optimization, i.e., learning the correspondence between hybrid visual features and labels, without considering the independent effects of states and objects. Thus, the model tends to capture correlations between states and objects that are preferred for the seen domain. For example, if the model has only seen sliced apple, it may tightly associate sliced with apple and fail to recognize sliced orange. An intuitive idea is to predict the labels of states and objects separately, excluding semantic interference with each other, which is the goal of decoupling representation. Decoupling representations strive to untangle states and objects, encouraging the model to extract semantic-independent visual representations. By recognizing states and objects individually, all seen or unseen compositions are in turn recognized. Decoupling representation is to learn the posteriori probabilities $p(s|\psi_{s}(x))$ and $p(o|\psi_{o}(x))$, where $\psi_{s}$ and $\psi_{o}$ are mapping functions. Mainstream approaches can be categorized as network design and attention separation.
IV-B1 Network Design
@@
Many researches resort to modifying the network structure to decompose visual features. Nan et al. [93] add two sub-networks at the bottom of the backbone network to extract state and object features, respectively, and then each predicts the corresponding labels and is trained with cross-entropy loss. CASUAL [42] employs a similar network structure. The difference is that it introduces a text-to-image mapping network that reverse maps text embeddings into visual space. Meanwhile, it uses the Hilbert-Schmidt Information Criterion, which is a non-parametric method for estimating the statistical dependence between samples of two random variables, to optimize training for ensuring state and object independence. ProtoProp [94] uses a prototypical network to learn independent representations. Specifically, regional features of images and prototypes of objects and states are computed for similarity and later optimized with cross-entropy loss. Hu et al. [95] propose a two-layer decoupling network. First, three sub-networks are used to extract state, object and composition features respectively. In the second layer, the state and object features are fed into a reorganization network to synthesize into a composition, while the composition feature is fed into two other sub-networks for further decoupling. CAILA [96] uses CLIP as a visual coder, which inserts three adapters in the middle of the network to capture the state, object, and composition semantic information, respectively. It is worth mentioning that the three-branch network has become a common infrastructure for most approaches, i.e., three sub-networks extract state, object, and composition visual features, respectively [97].
IV-B2 Attention Separation
@@
The motivation for attention separation is to dissect semantic features by capturing the parts of objects and states that are attended to separately. OADis [98] introduces a cross-sample attention mechanism to refine features. It equips each sample with two companion samples, one with the same object and different state, and the other with the same state and different object. Afterwards, it performs cross attention with the corresponding samples separately to refine the semantic features. ADE [99] adopts a similar training paradigm. Unlike pre-matched pairs, Jing et al. [100] utilize image retrieval techniques to find suitable matching samples for object and state branches, after which they are used to enhance the representations. Zhang et al. [101] propose a gradient-based attention decomposition method. They view compositional recognition as a domain generalization problem, and one of the solutions is to learn domain-invariant features. To this end, they measure the variation of features by evaluating the difference of the back-propagation gradients of sample pairs to recognize the domain-invariant part. DFSP [102] firstly decouples text embeddings by prototype computation, and after that, uses the text embeddings and image features for cross-attention in order to decompose the visual features. Similarly, Troika fuses multi-path prompt and cross attention to optimize text embeddings.
IV-C Dependency Modeling
@@
Decoupling techniques pursue the pure semantic representations, while mitigating the bias toward seen compositional concepts, introduce the bias toward single concept, i.e., state or object. This bias comes from the fact that the same concept is represented differently in diverse contexts. Due to the existence of complex acting mechanisms in the real world such as physical systems and biochemical reactions, the visual presentation of the same state on various objects varies greatly. For example, old is represented as metal rust spots on car, while it is shown as skin wrinkles on person. Obviously, concise text annotations cannot capture such subtle differences. And the independent visual representations learned by the model from the seen domain are biased, hindering the recognition accuracy on the unseen domain. To address this problem, many studies devote to modeling the dependency relationship between states and objects. By introducing a priori information about states or objects to correct each other’s semantic representations or classification confidence, they can be better generalized to the unseen domain. In short, the goal of dependency modeling is to learn the conditional probabilities $p(s|\psi_{s}(x),o)$ and $p(o|\psi_{o}(x),s)$. The relevant technologies according to the implementation objectives can be categorized into feature optimization and aid classification.
IV-C1 Feature Optimization
@@
The goal of feature optimization is to adjust each other’s feature representations by leveraging the semantic relevance of states and objects. Decoupled state/object features are expected to capture dependent semantic details while maintaining their own discriminative semantic information. Visual feature optimization is to correct the visual embedding distribution to avoid overfitting the seen domain caused by the model focusing on the single visual feature. For example, Panda et al. [104] set up multiple encoders between the state and object branching networks with the aim of transmitting object features to the state network in order to model the dependency of the state on the object. The same idea is taken by CoT [103], which uses object information as cueing signals while introducing an attention-based method to capture state-related visual features (Fig. 5). Another approach is to optimize text features. Since text annotations are usually pre-extracted, they do not have the ability to perceive dependency relationships. To this end, Nagarajan et al. [8] model states as operators, i.e., encode the state as a mapping function rather than an embedding, and thus explicitly learn the semantic embedding of an object ”after it has been operated on”. SymNet [105] introduces two sub-networks, the decoupling network and the coupling network, which are used to learn independent and compositional representations of objects and states. CANet [106] recognizes the object labels of an image firstly and then corrects the text embedding of the state by using the object text embedding as a priori condition.
IV-C2 Aid Classification
@@
Compared to adjusting the feature distribution, aid classification aims to directly introduce semantic information of objects/states to help the classifier to make discriminations. In earlier work, Chen et al. [2] utilize support vector machines (SVMs) for compositional recognition. They learn the weights of SVMs by capturing state dependencies on objects and generalize to unseen compositions. SAD-SP [107] takes an intuitive approach by introducing an attention module in addition to the conventional decoupling network. State features are used to predict object soft labels and vice versa. The predicted soft labels are then used to weight the predicted output of the conventional network. ProCC [108] performs information exchange at each layer of the state/object classifier. Specifically, it passes the output of one layer of the state classifier through a one-dimensional convolution and softmax, and then the output of the corresponding layer of the object classifier is multiplied with it and used as the input of the next layer. And vice versa.
IV-D Relationship Modeling
@@
Dependency modeling promotes the generalizability of models by exploring the intrinsic one-to-one state-object connections. Naturally, we wonder whether many-to-many complex mesh relationships, i.e., interactions between multiple objects or states, contribute to model performance. Such relationships describe the latent logical relations of elements in the real world due to various factors, such as subject affiliation, visual similarity, co-occurrence, etc. To name a few, we would spontaneously categorize broccoli and spinach, separating beef since they belong to different subjects; a fork and a spoon are more visually similar than an umbrella, and also they have a higher probability of co-occurrence. Intuitively, an understanding of this relations facilitates logical inference. It is because humans have a remarkable ability to draw analogies, make connections, and imagine, thus reasoning and recognizing by tapping into the relationships between different elements. When we see a furry black bear and know the similarities between polar bears and black bears, it is easy to migrate the state of furry to polar bears. Driven by this idea, relationship modeling aims to enhance the inference and representation capabilities of models by uncovering and maintaining potential relationships between states and objects. The dominant techniques can be summarized as cross attention, graph modeling and contrastive embedding.
IV-D1 Cross Attention

### v3

@@ 3.4 Attribute Component Analysis
Attributes serve as the crucial link between seen/unseen classes and provide essential supervisory signals for learning fine-grained visual representations. Consequently, the quality of the attribute space largely determines the upper limit of a recognition system’s performance. However, most attributes are manually created or pre-trained without proper adjustment, resulting in redundancy, noise, and unstructured relationships. To address this, attribute component analysis aims to optimize the attribute space through methods such as Attribute Enhancement and Refinement.
3.4.1 Attribute Enhancement
@@
Attribute enhancement aims to optimize attribute embeddings to improve their discriminability, using techniques like key attribute localization and attribute space optimization. Key attribute localization functions similarly to visual attention, identifying and focusing on the most critical parts. The cross attention mechanism mentioned in Equation 4 is one of the attribute localization techniques, just changing the input sequences. In addition, LFGAA [92] designs a latent guided attention module. It assumes that each image has different attention for attributes and, therefore, introduces multiple subnetworks to generate exclusive attribute weights for each image. Attribute space optimization adjusts pre-trained attribute embeddings to reduce semantic overlap that affects discrimination, especially in fine-grained datasets. For instance, many bird species share similar attributes, which makes distinguishing them challenging. For this reason, AREES [77] delineates the attribute embeddings with similar semantic distances into a cluster by K-means clustering algorithm, and the attributes in each cluster are subtracted from the cluster’s center of mass, i.e., the mean vector, to obtain the enhanced attribute vector. APNet [93] introduces graph modeling to discover the logical relationships between classes and updates the attributes by propagating the attribute features of neighboring nodes.
3.4.2 Attribute Refinement
@@
Since attributes are typically annotated by humans, the value they add from a human perspective to deep learning is uncertain. Attribute refinement seeks to address this by examining the design of attribute annotations, though few studies exist on the topic. MCZSL [94] points out that the quality of manual annotations is uncontrollable and time-consuming. Therefore, it attempts to design an automated attribute annotation scheme. Specifically, it obtains attribute descriptions from online media such as Wikipedia and designs a series of filtering strategies to filter out inappropriate attributes. DEDN [95] proposes a divide-and-conquer attribute learning scheme. It argues that attribute annotations are confusing and lead to semantic asymmetry problems. For example, for the attributes describing scenes, {studying, swimming} and {stone, tree} are not in the same dimension, with the former denoting function and the latter denoting object. Therefore, it designs a dual expert network in which one expert network focuses on processing attribute information in different dimensions.
4 Zero-Shot Compositional Recognition
4.1 Problem Formulation
@@
Suppose $\mathcal{S}=\{s_{1},s_{2},...,s_{n}\}$ is a set containing $n$ state elements and $\mathcal{O}=\{o_{1},o_{2},...,o_{m}\}$ is a set containing $m$ object elements. Then a compositional set can be obtained as $\mathcal{Y}=\{(s_{1},o_{1}),(s_{1},o_{2}),...,(s_{n},o_{m})|s\in\mathcal{S},o\in\mathcal{O}\}$, which containing $n\times m$ compositions. Now we have a seen domain $D^{s}=\{(x^{s},y^{s})|x^{s}\in\mathcal{X}^{s},y^{s}\in\mathcal{Y}^{s}\}$, where $\mathcal{X}^{s}$ denotes the training image set and $\mathcal{Y}^{s}\subseteq\mathcal{Y}$ is the label set. Similarly, we have an unseen domain denoted by $D^{u}=\{(x^{u},y^{u})|x^{u}\in\mathcal{X}^{u},y^{u}\in\mathcal{Y}^{u}\}$, where $\mathcal{X}^{u}$ denotes the test image set and $\mathcal{Y}^{u}\subseteq\mathcal{Y}$ is the label set. Data in the seen and unseen domains are disjoint, i.e., $X^{s}\cap X^{u}=\phi$ and $Y^{s}\cap Y^{u}=\phi$. The objective of ZSCR is to generalize to the unseen domain by learning the element-wise knowledge in the seen domain. Specifically, for closed-world compositional recognition, the goal is to learn a mapping function $\Psi:\mathcal{X}\to\mathcal{Y}^{s}\cup\mathcal{Y}^{u}$, wherein $\mathcal{X}^{u}$ is strictly inaccessible for training. Typically, the size of $\mathcal{Y}^{s}\cup\mathcal{Y}^{u}$ is smaller than that of $\mathcal{Y}$ since some compositions are not feasible. For open-world compositional recognition [27], the goal is to learn a mapping function $\Psi:\mathcal{X}\to\mathcal{Y}$, assuming that which compositions are infeasible is unknown.
@@
We divide the cutting-edge techniques of ZSCR into three categories based on their core components, including Decoupling Representation, Dependency Modeling, and Relationship Modeling. We then further split the sub-categories based on specific technical details or targets of action.
4.2 Decoupling Representation
@@
A common baseline approach involves mapping images to deep semantic space, computing similarity to text embeddings, and optimizing with cross-entropy loss. The goal is to learn the posterior probability $p(y|\psi(x))$, where $y=(s,o)$ and $\psi$ is a function mapping the image into the semantic space. This approach leads to model simplifying optimization by learning direct correspondences between visual features and labels, often missing independent effects of states and objects. For example, if the model has only seen sliced apple, it may tightly associate sliced with apple and fail to recognize sliced orange. An alternative is to separately predict the labels for states and objects to avoid semantic interference, known as decoupling representation. This approach aims to disentangle states and objects, encouraging the model to learn independent visual representations. By separately recognizing states and objects, the model can better generalize to new compositions. Decoupling representation focuses on learning posterior probabilities $p(s|\psi_{s}(x))$ and $p(o|\psi_{o}(x))$, where $\psi_{s}$ and $\psi_{o}$ are mapping functions. Mainstream approaches can be categorized as: Network Design and Attention Separation.
4.2.1 Network Design
@@
Many studies resort to modifying the network structure to decompose visual features. Nan et al. [96] add two sub-networks at the bottom of the backbone network to extract state and object features, respectively, and then each predicts the corresponding labels and is trained with cross-entropy loss. CASUAL [41] employs a similar network structure. The difference is that it introduces a text-to-image mapping network that reverse maps text embeddings into visual space. Meanwhile, it uses the Hilbert-Schmidt Information Criterion, which is a non-parametric method for estimating the statistical dependence between samples of two random variables, to optimize training for ensuring state and object independence. ProtoProp [97] uses a prototypical network to learn independent representations. Specifically, regional features of images and prototypes of objects and states are computed for similarity and later optimized with cross-entropy loss. Hu et al. [98] propose a two-layer decoupling network. First, three sub-networks are used to extract state, object, and composition features, respectively. In the second layer, the state and object features are fed into a reorganization network to synthesize into a composition, while the composition feature is fed into two other sub-networks for further decoupling. CAILA [99] uses CLIP as a visual coder, which inserts three adapters in the middle of the network to capture the state, object, and composition semantic information, respectively. It is worth mentioning that the three-branch network has become a common infrastructure for most approaches, i.e., three sub-networks extract state, object, and composition visual features, respectively [100].
4.2.2 Attention Separation
@@
The motivation for attention separation is to dissect semantic features by capturing the parts of objects and states attended to separately. OADis [101] introduces a cross-sample attention mechanism to refine features. It equips each sample with two companion samples, one with the same object and different state, and the other with the same state and different object. Afterwards, cross attention is performed with the corresponding samples separately to refine the semantic features. ADE [102] adopts a similar training paradigm. Unlike pre-matched pairs, Jing et al. [103] utilize image retrieval techniques to find suitable matching samples for object and state branches, after which they are used to enhance the representations. Zhang et al. [104] propose a gradient-based attention decomposition method. They view compositional recognition as a domain generalization problem, and one of the solutions is to learn domain-invariant features. To this end, they measure the variation of features by evaluating the difference in the back-propagation gradients of sample pairs to recognize the domain-invariant part. DFSP [105] firstly decouples text embeddings by prototype computation and, after that, uses the text embeddings and image features for cross-attention to decompose the visual features. Similarly, Troika [106] fuses multi-path prompt and cross attention to optimize text embeddings.
4.3 Dependency Modeling
@@
Decoupling techniques pursue the pure semantic representations while mitigating the bias toward seen compositional concepts, introducing the bias toward the single concept, i.e., state or object. This bias comes from the fact that the same concept is represented differently in diverse contexts. Due to the existence of complex acting mechanisms in the real world, such as physical systems and biochemical reactions, the visual presentation of the same state on various objects varies greatly. For example, old is represented as metal rust spots on car, while it is shown as skin wrinkles on person. Obviously, concise text annotations cannot capture such subtle differences, and the independent visual representations learned by the model from the seen domain are biased, hindering the recognition accuracy of the unseen domain. To address this problem, many studies are devoted to modeling the dependency relationship between states and objects. By introducing a priori information about states or objects to correct each other’s semantic representations or classification confidence, they can be better generalized to the unseen domain. In short, the goal of dependency modeling is to learn the conditional probabilities $p(s|\psi_{s}(x),o)$ and $p(o|\psi_{o}(x),s)$. The relevant technologies, according to the implementation objectives, can be categorized into Feature Optimization and Aid Classification.
4.3.1 Feature Optimization
@@
The goal of feature optimization is to adjust each other’s feature representations by leveraging the semantic relevance of states and objects. Decoupled state/object features are expected to capture dependent semantic details while maintaining their discriminative semantic information. Visual feature optimization corrects the visual embedding distribution to avoid overfitting the seen domain caused by the model focusing on a single visual feature. For example, Panda et al. [108] set up multiple encoders between the state and object branching networks to transmit object features to the state network to model the dependency of the state on the object. The same idea is taken by CoT [107], which uses object information as cueing signals while introducing an attention-based method to capture state-related visual features (Fig. 5). Another approach is to optimize text features. Since text annotations are usually pre-extracted, they can hardly perceive dependency relationships. To this end, Nagarajan et al. [6] model states as operators, i.e., encode the state as a mapping function rather than an embedding, and thus explicitly learn the semantic embedding of an object ”after it has been operated on”. SymNet [109] introduces two sub-networks, the decoupling network and the coupling network, which are used to learn independent and compositional representations of objects and states. CANet [110] recognizes the object labels of an image first and then corrects the text embedding of the state by using the object text embedding as a priori condition.
4.3.2 Aid Classification
@@
Compared to adjusting the feature distribution, aid classification aims to directly introduce semantic information of objects/states to make the classifier more discriminative. In earlier work, Chen et al. [2] utilize support vector machines (SVMs) for compositional recognition. They learn the weights of SVMs by capturing state dependencies on objects and generalize to unseen compositions. SAD-SP [111] takes an intuitive approach by introducing an attention module in addition to the conventional decoupling network. State features are used to predict object soft labels and vice versa. The predicted soft labels are then used to weigh the predicted output of the conventional network. ProCC [112] performs information exchange at each layer of the state/object classifier. Specifically, it passes the output of one layer of the state classifier through a one-dimensional convolution and softmax, and then the output of the corresponding layer of the object classifier is multiplied with it and used as the input of the next layer, and vice versa.
4.4 Relationship Modeling
@@
Dependency modeling promotes the generalizability of models by exploring the intrinsic one-to-one state-object connections. Naturally, we wonder whether many-to-many complex mesh relationships, i.e., interactions between multiple objects or states, contribute to model performance. Such relationships describe the latent logical relations of elements in the real world due to various factors, such as subject affiliation, visual similarity, co-occurrence, etc. To name a few, we would spontaneously categorize broccoli and spinach, separating beef since they belong to different subjects; a fork and a spoon are more visually similar than an umbrella, and also they have a higher probability of co-occurrence. Intuitively, an understanding of these relations facilitates logical inference. It is because humans have a remarkable ability to draw analogies, make connections, and imagine, thus reasoning and recognizing by tapping into the relationships between different elements. When we see a furry black bear and know the similarities between polar bears and black bears, it is easy to migrate the state of furry to polar bears. Driven by this idea, relationship modeling aims to enhance the inference and representation capabilities of models by uncovering and maintaining potential relationships between states and objects. The dominant techniques can be summarized as Cross Attention, Graph modeling, and Contrastive Embedding.
4.4.1 Cross Attention